INTRAMOLECULAR DYNAMICS OF THE N = 2 HF STRETCHING OVERTONE POLYAD OF (HF)_{2} STUDIED BY HIGH-RESOLUTION cw-DIODE LASER CAVITY RING-DOWN SPECTROSCOPY IN A PULSED SLIT JET

MICHAEL HIPPLER, \underline{LARS OELTJEN} and MARTIN QUACK, Physical Chemistry, ETH Zurich, CH-8093 Zurich, Switzerland.

The (HF)_{2} hydrogen bonded dimer has been a prototype system for high-resolution spectroscopy since the pioneering studies of its microwave spectra by Dyke, Howard, and Klemperer in 1972.^{1} Subsequently the HF stretching fundamentals were studied in 1983,^{2} a low frequency fundamental analyzed in the far infrared in 1987,^{3} HF stretching overtone spectra investigated by FTIR spectroscopy^{4} and finally full dimensional potential energy hypersurfaces developed of near to spectroscopic accuracy.^{5,6} All these were ”first” achievements prototypical for any type of hydrogen bonded dimer of this kind. Here we present the first study of the N = 2 HF stretching overtone polyad by very high resolution cw-diode laser cavity ring-down spectroscopy in pulsed slit jet expansions developed recently^{7} (instrumental bandwidth about 1 MHz corresponding to a resolving power of 2 × 10^{8}). An analysis of all polyad subbands in terms of spectroscopic constants, tunneling splittings, Lorentzian predissociation and Doppler contributions to the linewidths will be presented.^{8} The results agree well with full six-dimensional calculations^{9} but disagree with simple models or approximate calculations that have been presented in the past. 1 T. R. Dyke, B. J. Howard, and W. Klemperer, J. Chem. Phys. 56 (1972), 2442. 2 A. S. Pine, and W. J. Lafferty, J. Chem. Phys. 78 (1983), 2154. 3 K. von Puttkamer, and M. Quack, Mol. Phys. 62 (1987), 1047. 4 K. von Puttkamer, and M. Quack, Chem. Phys. 139 (1989), 31. 5 M. Quack, and M. A. Suhm, J. Chem. Phys. 95 (1991), 28. 6 W. Klopper, M. Quack, and M. A. Suhm, J. Chem. Phys. 108 (1998), 10096. 7 M. Hippler, and M. Quack, Chem. Phys. Lett. 314 (1999) 273; J. Chem. Phys. (2002) in press. 8 M. Hippler, L. Oeltjen, and M. Quack, in preparation. 9 J. Blumberger, L. Oeltjen, M. Quack, Z. Bac̆ić, and Y. Qiu, in preparation.